# REGEN: A Dataset and Benchmarks with Natural Language Critiques and Narratives

### v1

@@ 1. Introduction
This is challenging as existing datasets for training conversational recommender systems lack rich, contextual narratives and explicit user feedback. Such a dataset is crucial for modeling natural language interactions in the context of the recommended items, so generated text is consistent and relevant to the recommendations and user preferences. While several recent works explore modeling approaches that integrate the recommender’s internal state with an LLM (Zheng et al., 2024; Zhu et al., 2024; Hua et al., 2023; Ning et al., 2024b), they are limited to existing datasets (e.g. (Geng et al., 2022; Ni et al., 2019; Chen et al., 2024; Wang et al., 2024; Li et al., 2017; Liang et al., 2024)) that concentrate on user item-sequences, short dialog snippets, structured outputs, or small summaries that lack the rich and varied conversational element.
@@
Finally, to evaluate the utility of both the REGEN dataset and task, we present reference models and strong baselines. Here we introduce LUMEN (LLM-based Unified Multi-task Model with Critiques, Recommendations, and Narratives), a single model using an LLM backbone trained to to generate recommendations and narratives by integrating collaborative filtering signals, content signals, and natural-language critiques inputs. LUMEN leverages user-item interaction data, rich item content representations, and user direction to generate outputs that are informative and aligned with user preferences. Our experiments establish benchmarks that demonstrate the effectiveness of this architecture in producing human-like conversational recommendations.
@@ 3.1. Amazon Reviews
We use datasets from Amazon Product Reviews (Ni et al., 2019) as a basis for augmentation with rich narratives. The Amazon Reviews dataset (2018) is a massive collection of customer reviews spanning various product categories. This version, released in 2018, contains over 233 million reviews, making it one of the largest publicly available datasets for sentiment analysis and recommendation systems. Each review includes item features like title, description, category, price and review features including text, timestamp, score and summary. User sequences can be created by sorting each user review by timestamp, creating a sequential recommendation task. For our work, we use “Office Products” and “Clothing, Shoes and Jewelry,” verticals. These are chosen to be representative samples with different item counts, Office with  27k items and Clothing with  376k items. We plan to update these to other key verticals, and including the “All Reviews.”
@@ 3.2.1. Data Preprocessing:
User sequences $S_{u}=\{i_{1},i_{2},...,i_{n}\}$ consisting of interactions $i$ are first created by aggregating reviews per user, and sorting them by timestamp. For comparable results, we use the CALRec (Li et al., 2024) preprocessed datasets. Notably, CALRec’s de-duplication methodology helps mitigate trivial recommendation strategies, e.g. always recommending the most recent item. See Appendix A for the dataset statistics.
@@ 3.2.3. Critiques Generation
To generate critiques for adjacent item pairs that meet the similarity criteria, we query Gemini 1.5 Flash to synthesize several options, instructing the LLM to treat the first item as the current recommendation and the second item as representative of the user’s desired goal. The prompt contains the item descriptions and few shot examples. We select at random one of the generated critiques to inpaint into the dataset for that pair. For pairs that do not meet the similarity criteria, REGEN contains a sentinel placeholder. Dataset users can decide to train on the placeholder, to model the case where end-users do not provide critiques emulating a new search query, or to replace them with other critiques. For the “Clothing” dataset, REGEN includes LLM-generated critiques for about 18.6% of adjacent items appearing in user sequences. For “Offices,” there are about 17.6%.
@@ 3.2.4. Narrative Generation:
Using LLMs as automated evaluators, we assess generated critiques and narratives across multiple attributes. We demonstrate our approach using the Amazon product review dataset (Ni et al., 2019) and evaluate with Gemini Pro. Detailed descriptions and examples of the narratives are summarized in Appendix A, Table 5.
@@
Figure 2. Critiques evaluated by rated attributes i) Relevance ii) Specificity and iii) Quality for Office Products
@@ 4.2.2. LUMEN: Critiquing, Generative Recommendation and Narratives with LLM:
Table 2. Benchmarks with Joint Recommendation and Generation Tasks (Office Dataset)
@@
Hist. $\rightarrow$ ID + Narrative Recall@10 0.124 0.098 0.124 0.10 0.124 0.10 0.124 0.091 0.124 0.098
@@
Hist. + Critique $\rightarrow$ ID + Narrative Recall@10 0.1402 0.123 0.1402 0.117 0.1402 0.124 0.1402 0.122 0.1402 0.122
@@
Table 3. Benchmarks with Recommendation and Generation Tasks (Clothing Dataset) with Flare+LLM
Conversational Task Metrics
Purchase Long User
Reason Expl. Summary
Hist. $\rightarrow$ ID + Narrative Recall@10 0.1264
NDCG@10 0.1170
MRR 0.1152
BLEU 19.18 16.36
ROUGE 39.8 33.9
Sem. Sim. 0.892 0.921
Hist. + Critique $\rightarrow$ ID + Narrative Recall@10 0.1355
NDCG@10 0.1238
MRR 0.1215
BLEU 19.22 16.34
ROUGE 40.0 34.1
Sem. Sim. 0.892 0.921
@@ 5.1.2. Implementation Details
With the hybrid sequential recommender, we employed a Masked Language Modeling (MLM) task during training to enhance sample efficiency. We used the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$, and adopted the hyperparameter configurations detailed in (Hebert et al., 2024). Further, we made a simplifying assumption that the ground truth item is available to the LLM. Modeling the error in item recommendation, would require co-training or evaluating the sequential encoder and decoder LLM models together, which is difficult to reproduce. On the other hand, with this assumption, we see the results with this approach as an upper bound for the joint recommendation and generation tasks, that are easy to reproduce to benchmark the dataset, while also allowing for comparisons replacing the component models.
@@
5.1.3. Discussion of Results
The results are shown in Table 2 for the joint task of narratives. The “Flare + LLM” results, where recommendation and narrative generation are treated as separate tasks, provide an upper bound for the joint task performance. The recommendation metrics (Recall@10) for the joint task are within 10-15% of the metrics obtained with a sequential recommender. For the generative text metrics, the loss is primarily due to errors in decoding the item. Further, aggregate narratives such as user summaries, which depend more on the user than the item, show a smaller reduction in the joint task. In contrast, the purchase reason and product endorsement, which depend on an item context more than the user, show larger loss as expected, due to sensitivity to recommendation errors.
Table 3 shows results on the Clothing dataset, with a larger vocabulary size, using Flare+LLM. [^2] We will augment with the LUMEN results for the final revision of this paper. Compared to Office Products, we see similar scores on text metrics, with critiquing improving the recommendation metrics.
@@ 5.1.4. Validation with Model Traces
Appendix B presents model traces demonstrating natural language critiques effectively guiding recommendations and generating aligned narratives, showcasing the model’s responsiveness to critiques and its ability to produce preference-consistent narratives.
@@ 6. Limitations and Conclusion
This paper introduced REGEN, a novel dataset enhancing Amazon Reviews with personalized narratives and critiques, facilitating a new task in conversational recommendation. While a comprehensive human evaluation of REGEN’s critiques for “naturalness” would be beneficial, especially given their intended role as user feedback proxies, we argue that REGEN remains valuable even without it. Obtaining reliable human annotations for subjective, context-dependent textual tasks at scale is challenging. Moreover, the utility of a dataset is not solely dependent on perfect human alignment, as evidenced by the widespread use of Amazon Reviews despite its limitations. We demonstrate REGEN’s value by showing its effectiveness in developing conversational recommendation models that integrate sequential recommendations, critique generation, and narrative generation, further illustrated through traces from our benchmark models (Appendix B).
@@
Appendix A Additional Dataset Details
Table 4 show the dataset statistics. Details of the narratives and examples are summarized in Table 5.
Dataset # Users # Items I/U
Office 89,489 27,887 7.86
Clothing 1,168,735 376,847 9.21
Context Generated output Description Example
Appendix B Traces
Table 6 shows two recommender traces of LUMEN trained on the REGEN dataset.
Table 6. Model traces with the conversational task: with and without critiquing
Example 1
History: Sharpie 37003 Ultra-Fine Permanent Marker, Sharpie 37001 Permanent Markers, Sharpie Permanent Markers, Sharpie 37001 Permanent Markers
Target Item: Sharpie 15001 Box of 12 Sharpie Pro King Size Chisel Tip Permanent Markers
Critique: I need a bigger, chisel tip, for industrial use
Predicted Item (w/o Critique): Sharpie 37003 Ultra-Fine Permanent Marker
Predicted Item (w/ Critique): Sharpie 15001 Box of 12 Sharpie Pro King Size Chisel Tip Permanent Markers
Endorsement (w/o Critique): Upgrade your writing experience with the Sharpie 37001 Permanent Markers, Ultra Fine Point, Black, 12-pack! We see you’ve previously purchased and loved the Sharpie 12-pack of fine point markers, and this 12-pack of ultra-fine point markers offers the same high-quality, quick-drying, water-resistant, and fade-resistant ink in a convenient, easy-to-use format. Perfect for detailed work, note-taking, or any task requiring precision, these markers are ideal for both home and office use. Given your previous positive experiences with Sharpie markers, we’re confident you’ll find this 12-pack just as valuable and convenient.
Endorsement (w/ Critique): Based on your previous purchases of Sharpie Ultra Fine Point permanent markers in both black and purple, and your consistently high ratings, we believe you’ll love the Sharpie 37001 Industrial Chisel Tip Markers. These markers offer the same superior quality and long-lasting, fade-resistant ink you’ve come to expect from Sharpie, but with a larger chisel tip for bolder, more visible markings. Perfect for those larger-scale projects or those who prefer a more visible writing style, these markers are ideal for a variety of surfaces, including metal, plastic, and wood. Given your history of using Sharpie markers for various tasks, from labeling to artistic projects, these industrial-grade markers will provide the same reliable performance you’ve come to appreciate.”
Target Endorsement: Are you tired of markers that fade or don’t write on difficult surfaces? Based on your previous purchases of Sharpie Ultra Fine Point and Fine Point markers, we know you value quality, affordability, and a reliable writing tool. You’ve consistently given 5-star reviews, praising the sharp tips, long-lasting ink, and value for money compared to local stores. This Sharpie Pro King Size Permanent Marker builds on that foundation. Designed for heavy-duty use, it boasts an extra-large ink supply and a versatile chisel tip, perfect for writing on challenging surfaces like grease, wet or oily materials. This is the ideal choice for someone who, like you, needs a durable and high-performing marker for a variety of applications, and wants the same great value you appreciate in Sharpie products.”
Comments: This is a strong example because the critique (”bigger, chisel tip”) is clear and specific. The ”Predicted item w/ Critique” is exactly what the user asked for, while ”Predicted item w/o Critique” is the opposite. The endorsement with critique is also more compelling.
Example 2
History: Fellowes Booklift Copyholder, ACCO Binder Clips, Texas Instruments TI-84 Plus Graphics Calculator, Pentel Refill Ink for EnerGel 0.7mm Needle Tip Liquid Gel Pen
Target Item: Pentel Refill Ink for EnerGel Liquid Gel Pen, 0.5mm, Needle Tip, Red Ink
Critique: 0.5mm, not 0.7mm
Predicted Item (w/o Critique): Pentel Refill Ink for BL60 EnerGel Liquid Gel Pen, 1.0mm, Metal Tip, Red Ink
Predicted Item (w/ Critique): Pentel Refill Ink for EnerGel Liquid Gel Pen, 0.5mm, Needle Tip, Red Ink
Endorsement (w/o Critique): Upgrade your writing experience with the Pentel EnerGel Deluxe 0.7mm Retractable Liquid Gel Pen! We see you’ve previously enjoyed Pentel EnerGel pens, appreciating their smooth writing and vibrant colors. This deluxe model builds on that experience, offering a larger 0.7mm needle tip for even more precise writing and a comfortable, ergonomic design. The retractable tip and refillable design ensure long-lasting use, aligning with your previous purchases of high-quality writing instruments. This pen is perfect for detailed work, note-taking, or any task requiring precise and comfortable writing. Experience the difference of a premium writing instrument designed for both comfort and performance.
Endorsement (w/ Critique): Upgrade your writing experience with the Pentel EnerGel 0.5mm Retractable Gel Pens! Your previous purchases of Pentel EnerGel pens and refills show a clear preference for this high-quality, smooth-writing ink. You’ve consistently rated these pens highly, praising their vibrant colors and comfortable grip. This 0.5mm version offers the same superior ink quality in a smaller, more portable size, perfect for everyday use or note-taking. Given your history of purchasing multiple EnerGel pens and refills, this 0.5mm version is a natural progression, allowing you to enjoy the same exceptional writing experience in a more convenient format. Don’t miss out on this opportunity to experience the Pentel EnerGel difference!
Target Endorsement: Upgrade your writing experience with Pentel EnerGel 0.5mm Red Ink Refills! As a valued customer who previously purchased and highly rated both a copyholder and a set of mini binder clips, suggesting a focus on organization and efficiency in your workspace, these refills are designed to seamlessly integrate into your established system. Your positive feedback on both the TI-84 Plus graphing calculator and the Pentel EnerGel 0.7mm refills shows a preference for high-quality, reliable writing instruments. These 0.5mm refills offer the same smooth, quick-drying ink you loved, now in a slightly finer point perfect for detailed note-taking or precise writing. A 12-pack provides long-lasting convenience, ensuring you always have the right ink on hand to keep your work flowing smoothly – just like the superior organization you’ve shown with your previous purchases. Get the Pentel EnerGel 0.5mm Red Ink Refills today and experience the same smooth, reliable quality you’ve come to expect from Pentel, without interruption.
Comments: The critique (”0.5mm, not 0.7mm”) is clear and specific. The ”Predicted item w/ Critique” is precisely what the user asked for. The endorsement with critique directly addresses the user’s specific need.

### v2

@@ 1. Introduction
This is challenging as existing datasets for training conversational recommender systems lack rich, contextual narratives and explicit user feedback. Such a dataset is crucial for modeling natural language interactions in the context of the recommended items, so generated text is consistent and relevant to the recommendations and user preferences. While several recent works explore modeling approaches that integrate the recommender’s internal state with an LLM (Zheng et al., 2024; Zhu et al., 2024; Hua et al., 2023; Ning et al., 2024b), they are limited to existing datasets (e.g. (Geng et al., 2022; Ni et al., 2019; Chen et al., 2024; Wang et al., 2024; Li et al., 2017; Liang et al., 2024)) that concentrate on user item-sequences, short dialog snippets, structured outputs, or small summaries that lack the rich and varied conversational elements.
@@
Finally, to evaluate the utility of both the REGEN dataset and task, we present reference models and strong baselines. Here we introduce LUMEN (LLM-based Unified Multi-task Model with Critiques, Recommendations, and Narratives), a single model using an LLM backbone trained to to generate recommendations and narratives by integrating collaborative filtering signals, content signals, and natural-language critiques inputs. LUMEN leverages user-item interaction data, rich item content representations, and user feedback to generate outputs that are informative and aligned with user preferences. Our experiments establish benchmarks that demonstrate the effectiveness of this architecture in producing human-like conversational recommendations.
@@ 3.1. Amazon Reviews
We use datasets from Amazon Product Reviews (Ni et al., 2019) as a basis for augmentation with rich narratives. The Amazon Reviews dataset (2018) is a massive collection of customer reviews spanning various product categories. This version, released in 2018, contains over 233 million reviews, making it one of the largest publicly available datasets for sentiment analysis and recommendation systems. Each review includes item features like title, description, category, price and review features including text, timestamp, score and summary. User sequences can be created by sorting each user review by timestamp, creating a sequential recommendation task. For our work, we use “Office Products” and “Clothing, Shoes and Jewelry,” verticals. These are chosen to be representative samples with different item counts, Office with  27k items and Clothing with  376k items.
Dataset # Users # Items I/U
Office 89,489 27,887 7.86
Clothing 1,168,735 376,847 9.21
@@ 3.2.1. Data Preprocessing:
User sequences $S_{u}=\{i_{1},i_{2},...,i_{n}\}$ consisting of interactions $i$ are first created by aggregating reviews per user, and sorting them by timestamp. For comparable results, we use the CALRec (Li et al., 2024) preprocessed datasets. Notably, CALRec’s de-duplication methodology helps mitigate trivial recommendation strategies, e.g. always recommending the most recent item. See Table 1 for the dataset statistics.
@@ 3.2.3. Critiques Generation
To generate critiques for adjacent item pairs that meet the similarity criteria, we query Gemini 1.5 Flash to synthesize several options, instructing the LLM to treat the first item as the current recommendation and the second item as representative of the user’s desired goal. The prompt contains the item descriptions and few shot examples. We select at random one of the generated critiques to inpaint into the dataset for that pair. For pairs that do not meet the similarity criteria, REGEN contains a sentinel placeholder. Dataset users can decide to train on the placeholder, to model the case where end-users do not provide critiques emulating a new search query, or to replace them with other critiques. For the “Clothing” dataset, REGEN includes LLM-generated critiques for about 18.6% of adjacent items appearing in user sequences. For “Office” dataset there are about 17.6%.
@@ 3.2.4. Narrative Generation:
Using LLMs as automated evaluators, we assess generated critiques and narratives across multiple attributes. We demonstrate our approach using the Amazon product review dataset (Ni et al., 2019) and evaluate with Gemini Pro. Detailed descriptions and examples of the narratives are summarized in Table 2.
Context Generated output Description Example
@@
Figure 3. Critiques evaluated for i) Relevance ii) Specificity and iii) Quality for Office Products
@@ 4.2.2. LUMEN: Critiquing, Generative Recommendation and Narratives with LLM:
Table 4. Office Products Dataset - Benchmarks with Joint Recommendation and Generation Tasks. To evaluate upper bound performance, we provide the true next item $\mathrm{ID}_{t+1}$ when generating Flare+LLM narratives.
@@
Hist. $\downarrow$ ID + Narrative Recall@10 0.124 0.098 0.124 0.10 0.124 0.10 0.124 0.091 0.124 0.098
@@
Hist. + Critique $\downarrow$ ID + Narrative Recall@10 0.1402 0.123 0.1402 0.117 0.1402 0.124 0.1402 0.122 0.1402 0.122
@@
Table 5. Clothing Dataset - Benchmarks with Joint Recommendation and Generation Tasks. To evaluate upper bound performance, we provide the true next item $\mathrm{ID}_{t+1}$ when generating Flare+LLM narratives.
Purchase Reason User Summary
Task Metrics Flare+LLM LUMEN Flare+LLM LUMEN
Hist. $\downarrow$ ID + Narrative Recall@10 0.126 0.127 0.126 0.118
NDCG@10 0.117 0.111 0.117 0.100
MRR 0.115 0.106 0.115 0.094
BLEU 19.2 10.2 16.4 12.8
ROUGE 39.8 29.7 33.9 29.9
Sem. Sim. 0.89 0.78 0.92 0.90
Hist. + Critique $\downarrow$ ID + Narrative Recall@10 0.136 0.143 0.136 0.151
NDCG@10 0.124 0.122 0.124 0.134
MRR 0.122 0.115 0.122 0.128
BLEU 19.2 10.5 16.3 13.3
ROUGE 40.0 30.2 34.1 30.3
Sem. Sim. 0.89 0.79 0.92 0.90
@@ 5.1.2. Implementation Details
With the hybrid sequential recommender, we employed a Masked Language Modeling (MLM) task during training to enhance sample efficiency. We used the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$, and adopted the hyperparameter configurations detailed in (Hebert et al., 2024). Further, we made a simplifying assumption that the ground truth item is available to the LLM. Modeling the error in item recommendation would require co-training or evaluating the sequential encoder and decoder LLM models together, which is difficult to reproduce. On the other hand, with this assumption, we view the results with this approach as an upper bound for the joint recommendation and generation tasks. The approach is also easily reproducible, allowing for comparisons of other component models.
@@
Figure 6. An example trace from LUMEN, using a joint generative task on REGEN, demonstrates the conversational recommender model’s ability to process purchase interactions and in-context natural language critiques. This processing improves the predicted item and generates a natural language endorsement that aligns with the prediction.
5.2. Results and Discussion
Our experimental results, summarized in Table 4 and Table 5, lead to the following key findings:
5.2.1. Recommendation Performance
To establish baselines, we first compare recommendation metrics, which are well-studied for several recommender models in related work on amazon datasets.
On REGEN Clothing (Table 5), compared to FLARE (See (Hebert et al., 2024) and references within) as a baseline, LUMEN trained on joint task achieves competitive performance despite generating recommendations and narratives within a single model, often achieving within ±$\pm 0.01$ Recall@10 and MRR, and improving performance in some cases.
For Office Products (Table 4) however, we see slightly lower recommendation metrics, within 10-15% of the baseline. The Office dataset has a significantly smaller vocabulary ($\approx$27k) and training data size ($\approx$90k) (Table 1). We believe this indicates that a sufficiently large training dataset is required to retain the recommendation performance in models trained on joint tasks.
5.2.2. Recommendation Performance with Natural Language Critiques
Incorporating user critiques consistently enhances recommendation performance, with improvements of +0.01–0.02 in Recall@10 and NDCG@10 and +0.01–0.03 in MRR observed for both hybrid and end-to-end LUMEN models ( Table 5).
5.2.3. Narrative Performance
As expected, LUMEN’s narrative quality scores (BLEU, ROUGE) are lower than those of a hybrid model conditioned directly on ground-truth items. We attribute this performance gap primarily to recommendation errors, as LUMEN’s narratives are based on its own predicted items in a single end to end model. We find that this dependency was less impactful for aggregate narratives such as “User Summary,” which rely less on the accuracy of the most recent item recommendation and more on the overall user history. In this case, LUMEN compares well with an LLM trained to generate narratives alone. Further, we expect that an integrated hybrid architecture where the output of the recommender is directly used as input to LLM is likely to see similar errors.
@@ 5.2.4. Validation with Model Traces
Figure 6 shows an example trace demonstrating recommendation, natural language understanding and generation capabilities of the LUMEN model trained on the joint tasks with the REGEN dataset.
@@ 6. Limitations and Conclusion
This paper introduced REGEN, a novel dataset enhancing Amazon Reviews with personalized narratives and critiques, facilitating a new task in conversational recommendation. While a comprehensive human evaluation of REGEN’s critiques for “naturalness” would be beneficial, especially given their intended role as user feedback proxies, we argue that REGEN remains valuable even without it. Obtaining reliable human annotations for subjective, context-dependent textual tasks at scale is challenging. Moreover, the utility of a dataset is not solely dependent on perfect human alignment, as evidenced by the widespread use of Amazon Reviews despite its limitations. We demonstrate REGEN’s value by showing its effectiveness in developing conversational recommendation models that integrate sequential recommendations, critique generation, and narrative generation.